INTERSTELLAR UREA [(NH_{2})_{2}CO]

\underline{LEWIS E. SNYDER}, ANTHONY J. REMIJAN, FRANCIS J. LOVAS^{a}, DOUGLAS N. FRIEDEL, Department of Astronomy, The University of Illinois, 1002 W. Green, Urbana, IL 61801; SHENG-YUAN LIU^{b}, Astronomy Department, California Institute of Technology, Pasadena, CA 91125.

In 1773, H. M. Rouelle’s discovery of urea in human urine supported theories that organic chemicals could be modified by chemistry, but could only be produced through a vital force in living plants and animals^{a}. The theory of Vitalism was badly damaged in 1828 when Friedrich Wöhler synthesized urea from inorganic starting materials. Indeed, urea was the first organic compound ever to be synthesized from inorganic compounds. Today, urea is thought to have played a significant role in prebiotic chemistry. Extraterrestrial urea was first detected in two samples of the Murchinson meteorite^{b}. In this paper, we report the first detection of gas-phase interstellar urea [(NH_{2})_{2}CO]. We used the BIMA^{c} Array to detect urea emission from the hot molecular core Sgr B2(N-LMH) at λ = 3 mm (9_{∗,9}-8_{∗,8}, 9_{∗,8}-8_{∗,7}, and 10_{∗,10}-9_{∗,9}), and at λ = 1 mm (20_{∗,20}-19_{∗,19}, 20_{∗,19}-19_{∗,18}, and 21_{∗,21}-20_{∗,20}). We will compare the urea column density and distribution with those of other molecules of biological interest in Sgr B2(N-LMH).

aCurrent Address: Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441

bCurrent address: Institute of Astronomy and Astrophysics, Academia Sinica, P. O. Box 23-141,106 Taipei, Taiwan, ROC

aFairall 1996, http://smallfry.dmu.ac.uk/chem/mom/urea/urea.html

bHayatsu et al. 1975, Geochim. et Cosmochim. Acta, 39, 471

cOperated by the University of California, Berkeley, the University of Illinois, and the University of Maryland with support from the National Science Foundation